OpenStudy (anonymous):

Use the zeros of the following quadratic to find the x-value of the vertex. 2003-18-02-00-00_files/i0170000.jpg

OpenStudy (anonymous):

y=x^2-2x+1

OpenStudy (anonymous):

where the image

OpenStudy (anonymous):

its that problem i just posted

OpenStudy (anonymous):

oh ok... one sec

OpenStudy (anonymous):

y=x^2-2x+1 y-1=x^2-2x y-1/2=x^2-x y-1/2=x^3

OpenStudy (anonymous):

those answers arent showing up

OpenStudy (anonymous):

a.1 b.-1 c.-2 d.2

OpenStudy (anonymous):

y=x^2-2x+1 y=1^2-2(1)+1 y=1-2+1 y=0

OpenStudy (anonymous):

just solve the rest just like that

OpenStudy (anonymous):

thanks!!!!!